OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? five, six, eleven five, nine, ten five, thirteen, eighteen five, twelve, thirteen

jimthompson5910 (jim_thompson5910):

Let's say we have a triangle with sides of a,b,c where c is the longest side. If a^2 + b^2 = c^2 is true, then we have a right triangle

jimthompson5910 (jim_thompson5910):

For example, if you have a triangle with sides 3,4,5, then we have a^2 + b^2 = c^2 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25 that shows us the triangle with sides 3,4,5 is a right triangle